# The Fall of an Algorithm: Characterizing the Dynamics Toward Abandonment

### v1

@@ Abstract.
As more algorithmic systems have come under scrutiny for their potential to inflict societal harms, an increasing number of organizations that hold power over harmful algorithms have chosen (or were required under the law) to abandon them. While social movements and calls to abandon harmful algorithms have emerged across application domains, little academic attention has been paid to studying abandonment as a means to mitigate algorithmic harms. In this paper, we take a first step towards conceptualizing “algorithm abandonment” as an organization’s decision to stop designing, developing, or using an algorithmic system due to its (potential) harms. We conduct a thematic analysis of real-world cases of algorithm abandonment to characterize the dynamics leading to this outcome. Our analysis of 40 cases reveals that campaigns to abandon an algorithm follow a common process of six iterative phases: discovery, diagnosis, dissemination, dialogue, decision, and death, which we term the 6 D’s of abandonment. In addition, we highlight key factors that facilitate (or prohibit) abandonment, which include characteristics of both the technical and social systems that the algorithm is embedded within. We discuss implications for several stakeholders, including proprietors and technologists who have the power to influence an algorithm’s (dis)continued use, FAccT researchers, and policymakers.
@@ 1. Introduction
00footnotetext: † Both senior authors contributed equally to this research.An increasing number of incidents have demonstrated how algorithmic systems can inflict societal harm, either by upholding oppressive power structures (Noble, 2018; Benjamin, 2019), reinforcing stereotypes (Vincent, [n. d.]; Andalibi et al., 2023), or making life-altering decisions through inscrutable logic (Selbst and Barocas, 2018; Liban, 2018; Forrest, 2021; Eppink, 2023). In response to these harms, a growing body of research has explored how harmful algorithms might be repaired through technical or procedural interventions (Schwennesen, 2019; Bembeneck et al., 2021; Orphanou et al., 2022). However, scholars and activists alike have argued that not all algorithmic harms can be effectively addressed or mitigated through these kinds of repairs (Harcourt, 2006; Milner, 2020; Wang et al., 2023). This stance often arises from a recognition of the deeply ingrained nature of the corresponding harm, where mere adjustments fail to resolve fundamentally systemic issues. In some instances, resistance from impacted communities has precipitated the fall of an algorithmic system, in which its proprietors might choose (or be forced to) abandon its use. Take, for example, a case from Arkansas: when a new resource allocation algorithm resulted in a drastic cut to the number of in-home care hours that MedicAid beneficiaries would receive (Hub, 2022a), one beneficiary wrote in with a clear request: to turn it off, and return to the old system of human-based assessment. No one could offer an explanation for why this beneficiary’s in-home care hours were reduced, cuts that for many beneficiaries with chronic disabilities resulted in drastic changes to their quality of life. While three years later, the algorithm (called RUGS) was abandoned, the road to this outcome was long and tedious. Efforts to abandon the algorithm were met with resistance from state officials (Weil and Edwards, 2023). The eventual decommissioning of RUGS after it was invalidated by a state court (Hub, 2022a) was a result of concerted advocacy efforts by impacted individuals, legal advocates, and media attention.
@@
While some campaigns to abandon harmful algorithms have been successful, other harmful algorithms continue to remain in use today. Furthermore, while some algorithms were abandoned in early ideation phases before they were deployed and some were abandoned within mere hours of harm discovery, other algorithms continue(d) to inflict harm for years. Thus, open questions remain when considering the possibility of abandoning harmful algorithms as a mechanism of harm mitigation and accountability: What types of events preceded and precipitated the decision to abandon an algorithm? How did impacted communities and other actors work together to contest harmful algorithms? What technical, political, or social characteristics of real-world algorithmic deployments posed barriers for contestation, accountability, and abandonment? How can future calls for abandonment be bolstered and supported?
@@
In this paper, we take a first step toward answering these questions by exploring the concept of algorithm abandonment. We define algorithm abandonment as a decision made by actors with jurisdiction over the system, to discontinue the process of developing, deploying, or using the algorithm due to its (potential) harms. The decision to abandon an algorithm can be made at any point in the algorithms’ development lifecycle (Hub, 2022d; Suresh and Guttag, 2021), and includes both decisions to not build a proposed technology and decisions to stop using tools that are already built. The algorithmic systems that we explore span diverse ways of coming into being (Wang et al., 2023), including hand-coded rules and machine learning models. Our study of algorithm abandonment draws from literature from Science and Technology Studies which examines algorithms not just as technical artifacts mapping inputs to outputs, but rather as socio-technical “assemblages” of social relationships that enable and sustain technical artifacts (Schwennesen, 2019). Thus, understanding abandonment requires examining the social, political, and economic systems that algorithms are embedded within (Seaver, 2019).
@@
In our thematic analysis of 40 different campaigns to abandon harmful algorithms in recent years, we observe that the fall of an algorithm (or “dynamics” leading to abandonment) follows a consistent pattern of six iterative, non-linear phases which we term the 6 “D”s of abandonment: discovery, diagnosis, dissemination, dialogue, decision, and death, which we visualize for three example cases in Figure 1. We review each of these phases and highlight both tactics that algorithm owners used to avoid scrutiny (e.g., discrediting or ignoring criticism) and tactics that other actors such as impacted individuals, advocacy organizations, journalists, or regulators, used to shift power to advocates of abandonment. We additionally identify key socio-technical factors most relevant to the outcome of whether an algorithm was ultimately abandoned (and how long it took to reach this stage), such as public awareness of and access to the algorithm, technicalities of the algorithm’s implementation, visibility of criticism, the regulatory environment, and for whom (and how) the algorithm is marketed. We conclude by discussing implications for technologists designing algorithmic systems, regulators and oversight entities, and the FAccT community in identifying barriers to abandonment and interventions to support future calls to abandon harmful algorithms. We release our annotated case set online111https://njohnson99.github.io/fall-of-algorithm-database/ as a resource to aid future research and education.
@@ 2. Background & Related Work
We contextualize our study of abandonment using scholarship that studies how different stakeholders respond to algorithmic harm. “Algorithmic harm,” as described by Shelby et al. (2023), refers to “the adverse lived experiences resulting from a system’s deployment and operation in the world” which occurs through the interplay of technical components and societal power dynamics (Noble, 2018; Benjamin, 2019; McGregor, 2020). We group approaches taken to address algorithmic harm into two broad categories: (1) actions taken “from above” by organizations in positions of power (e.g., those that develop, oversee, or govern the algorithm’s use), and (2) actions taken “from below” by other stakeholders (e.g., those impacted by algorithmic harm or the public). We adopt the language of “above” and “below” from Whitney et al. (2021), who uses these terms to articulate the “tensions between the wants, needs, and knowledge of unequally empowered stakeholders”.
@@
We situate our study of abandonment as an organizational decision made “from above” by algorithm owners or government oversight entities. We use the term “algorithm owner” to broadly refer to the organizations and people within them that have the power to influence an algorithm’s (dis)continued use on the ground. Our analysis of the events, or “dynamics”, that precede abandonment in practice draws from the growing body of scholarship that examines how actors “from below” organize to contest harmful algorithmic systems (Zong and Matias, 2023; DeVrio et al., 2023). We use the term “impacted communities” to describe the people who are affected by the use of an algorithm. Those impacted may include the consumers of a technology, or the people who are subject to an algorithm’s decisions (Friedman et al., 2006; Moss et al., 2021).
@@
Repair from above. Much work concerning algorithmic harm explores the actions that algorithm owners might take to repair flawed technologies. The AI/ML and FAccT community have proposed numerous technical methods to promote values of “fairness” and “explainability” (Doshi-Velez and Kim, 2017; Selbst and Barocas, 2018; Liao and Vaughan, 2023). This work has led to the development of metrics (Speicher et al., 2018; Pessach and Shmueli, 2022; Barocas et al., 2023; Subhash et al., 2022), methods (Dwork et al., 2012; Wang and Rudin, 2015; Hardt et al., 2016; Ribeiro et al., 2016), and and toolkits (Bellamy et al., 2018; Arya et al., 2019; Deng et al., 2022; Weerts et al., 2023) that practitioners can use to measure and “fix” specific harmful behaviors. Beyond technical repairs, human-computer interaction researchers have documented how organizational users and decision-makers engage in their own ad-hoc or improvisational repair activities when they encounter flawed technologies (Houston et al., 2016; Eslami et al., 2017; Schwennesen, 2019). For example, social workers using AI tools to screen calls can override them to mitigate disparities in final outcomes (De-Arteaga et al., 2020; Cheng and Chouldechova, 2022; Cheng et al., 2022). Such methods can be used to correct the algorithm’s flaws, often defined in terms of narrowly scoped measures (Jacobs and Wallach, 2021), while maintaining the algorithm’s operation.
@@
Refusal from below. The concept of refusal, defined simply as the act of saying “no” (Ahmed, 2017), has been studied in feminist and de-colonial discourses (Simpson, 2007; TallBear, 2013; McGranahan, 2016; Honig, 2021). Today, scholars and activists continue to draw from these traditions in studying refusal of new technologies (Cifor et al., 2019; DeVrio et al., 2023; Zong and Matias, 2023). Among theories of anti-techno-solutionism, Benjamin (2016) refers to indigenous refusal to analyze vignettes where research subjects refuse to engage with new biotechnologies. A group of scholars drafted the Feminist Data “Manifest-No”, a declaration of refusal to harmful data regimes and exploitative research practices (Cifor et al., 2019). There is a rich history of grassroots social movements formed to refuse technologies, tracing back hundreds of years to the British Luddites during the Industrial Revolution (Conniff, 2011), sparking modern adaptations of “Luddism” by workers organizing to oppose harmful algorithmic management practices (Chayka, 2023; Merchant, 2023).
Refusal of harmful technologies may take the form of individual users choosing to no longer interact with harmful algorithms (DeVrio et al., 2023; Zong and Matias, 2023). For example, in response to Amazon’s introduction of AI-powered cameras for worker surveillance, some drivers engaged in an act of refusal by obscuring the cameras with stickers, rendering them unable to record (Gurley, 2021). But refusal also extends beyond individual acts, into collective calls for a system to be abandoned. Understanding calls to abandon an algorithm through the lens of refusal sheds light on the possibilities that abandonment entails to “open up possibilities for power shifts” (Casemajor et al., 2015).
@@ 3. Methodology
Case collection. To develop our dataset, we conducted an extensive search of various sources for instances where an algorithm that caused harm was either (a) abandoned, or (b) critics called for the algorithm to be abandoned, regardless of whether it was abandoned or not. Our search methodology was a dynamic, iterative process that unfolded concurrently with our study. This approach involved consulting a diverse group of resources, detailed further below:
@@
AI incident databases: We reviewed all of the reported incidents to date222We conducted our review from July to December 2023, and reviewed all incidents that had been posted in each database before January 1, 2024. in three incident databases that compiled and documented real-world harms caused by algorithmic systems: (a) the “AI Incident Database” (AIID) (McGregor, 2020), (b) the “AI, Algorithmic, and Automation Incidents and Controversies” (AIAAIC) repository (aia, [n. d.]), and (c) “Not My AI” (Varon and Peña, 2022) which focuses on mapping harmful algorithms in Latin America.
@@
International cases: Given that U.S. and European-based incidents were over-represented in our initial case set, we extended our search to be intentionally inclusive of other regions globally. We conducted additional keyword searches on Google (detailed in Appendix B) to identify relevant cases that occurred in countries outside of Western contexts, such as Korea (Kwon and Yun, 2021), Argentina (Aplicada, 2018), Jordan (Ryan-Mosley, 2023a), and India (amn, 2022).
@@
| U.S. National Eating Disorder Assocation (NEDA)’s Tessa Wellness Bot (Wells, 2023) (chatbot) | Shortly after NEDA fired their hotline staff, eating disorder activists posted on social media about how Tessa (NEDA’s chatbot) gave harmful and triggering advice. |
@@
From the cases that met the abandonment criteria, we selected a subset of 40 as our case studies for our thematic analysis (which we describe in the following section). We chose this subset based on several criteria, including a) coverage of different domains, b) access to enough sources to gain more information about a case to do a thematic analysis, and c) reaching saturation in finding new themes in the analysis process. The selected cases included algorithms used in social media, chatbots, hiring, family and social services, benefits allocation, healthcare, education, property tech, and law enforcement. We chose among cases within the same domain based on the amount of public information and documentation that was available for us to construct accurate event timelines. We performed our thematic analysis in parallel to identifying relevant cases, and stopped adding new cases to the subset once we had reached saturation in finding new themes (i.e., the types of events and factors relevant to abandonment). Our final case set contains 40 different algorithmic systems, 24 of which had been successfully abandoned at the time of writing333Some algorithms that we study are deployed in multiple different contexts (e.g., PredPol was used by dozens of U.S. police departments (Haskins, 2019)), and were abandoned in some, but not others. We ‘count’ products like PredPol as a single “algorithmic system”, and count them as being “abandoned” if they were abandoned in at least one such context. , while the remainder were still in use. 16 out of the 40 total cases involved campaigns that took place in countries outside of the U.S. To provide context, we describe three example cases in Table 1 and a complete list in Appendix A. These 40 cases are not intended to be a comprehensive list of all algorithms that were called to be abandoned, but rather serve as a starting point to conceptualize algorithm abandonment, illustrate its dynamics, and map out its implications.
@@
Thematic analysis. We gathered and reviewed public resources about each case, including news articles, research publications, social media posts, and any other relevant primary sources (e.g., algorithm source code or court documents). For each case, we constructed a timeline by noting the month and year of important events. Our goal in noting important events was not to be comprehensive, but rather to provide “good enough” coverage of the types of events leading to abandonment, a standard used in past qualitative research (Yildirim et al., 2023). We reviewed sources to code metadata about the algorithm, such as its application domain, the type(s) of algorithmic harm (Shelby et al., 2023), and other information relevant to whether it was ultimately abandoned (detailed in Appendix B).
@@
We conducted a bottom-up thematic analysis (Braun and Clarke, 2006) of case event timelines and metadata in a series of meetings among research team members. In our thematic analysis of event timelines, we paid particular attention to the sequencing of key events, the timelapse between them, and how each event contributed to the final decision of whether the system was abandoned. To identify factors relevant to abandonment presented in Section 5, we noted differences among algorithms that were abandoned almost immediately (e.g., within hours of their release), versus algorithms that took months or even years to be abandoned. We organized these differences into higher-level categories that resulted in the final seven factors. We also noted the causes for (call for) abandonment, different barriers that advocates of abandonment faced when asserting their role in this accountability relationship, and the broader implications for research on algorithmic accountability. We walk through the set of sources and resulting codes and event timelines for an example case in Appendix B.1 . Our thematic analysis yielded six high-level types of events (or “phases”) as the dynamics of the algorithm abandonment that we describe in Section 4.

### v2

@@ Abstract.
As more algorithmic systems have come under scrutiny for their potential to inflict societal harms, an increasing number of organizations that hold power over harmful algorithms have chosen, or were required under the law, to abandon them. While social movements and calls to abandon harmful algorithms have emerged across application domains, little academic attention has been paid to studying abandonment as a means to mitigate algorithmic harms. In this paper, we take a first step towards conceptualizing “algorithm abandonment” as an organization’s decision to stop designing, developing, or using an algorithmic system due to its (potential) harms. We conduct a thematic analysis of real-world cases of algorithm abandonment to characterize the dynamics leading to this outcome. Our analysis of 40 cases reveals that campaigns to abandon an algorithm follow a common process of six iterative phases: discovery, diagnosis, dissemination, dialogue, decision, and death, which we term the 6 D’s of abandonment. In addition, we highlight key factors that facilitate (or prohibit) abandonment, which include characteristics of both the technical and social systems that the algorithm is embedded within. We discuss implications for several stakeholders, including proprietors and technologists who have the power to influence an algorithm’s (dis)continued use, FAccT researchers, and policymakers.
@@ 1. Introduction
An increasing number of incidents have demonstrated how algorithmic systems can inflict harm, by (for example) upholding oppressive power structures (Noble, 2018; Benjamin, 2019), reinforcing stereotypes (Vincent, [n. d.]; Andalibi et al., 2023), or making life-altering decisions through inscrutable logic (Selbst and Barocas, 2018; Liban, 2018; Forrest, 2021; Eppink, 2023). In response to these harms, a growing body of research has explored how harmful algorithms might be repaired through technical or procedural interventions (Schwennesen, 2019; Bembeneck et al., 2021; Orphanou et al., 2022). However, scholars and activists alike have argued that not all algorithmic harms can be effectively addressed or mitigated through these kinds of repairs (Harcourt, 2006; Milner, 2020; Wang et al., 2023). This stance often arises from a recognition of the deeply ingrained nature of the corresponding harm, where mere adjustments fail to resolve fundamentally systemic issues. In some instances, resistance from impacted communities has precipitated the fall of an algorithmic system, in which its proprietors might choose, or be forced to, abandon its use.
Take, for example, a case from Arkansas: when a new resource allocation algorithm resulted in significant cuts to the number of in-home care hours that Medicaid beneficiaries would receive (Hub, 2022a), one beneficiary wrote in with a clear request: to turn it off, and return to the old system of human-based assessment. No one in the state could offer an explanation for why this beneficiary’s in-home care hours were reduced, cuts that for many beneficiaries with chronic disabilities resulted in drastic changes to their quality of life. While three years later, the algorithm (called RUGS) was abandoned, the road to this outcome was long and tedious. Efforts to abandon the algorithm were met with resistance from state officials (Weil and Edwards, 2023). The eventual decommissioning of RUGS after it was invalidated by a state court (Hub, 2022a) was a result of concerted advocacy efforts by impacted individuals, legal advocates, and media attention.
@@
Although some campaigns to abandon harmful algorithms have been successful, other harmful algorithms continue to remain in use today. Furthermore, while some algorithms were abandoned in early ideation phases before they were deployed, and some were abandoned within mere hours of harm discovery, other algorithms continue(d) to inflict harm for years. Thus, open questions remain when considering the possibility of abandoning harmful algorithms as a mechanism of harm mitigation and accountability: What types of events preceded and precipitated the decision to abandon an algorithm? How did impacted communities and other actors work together to contest harmful algorithms? What technical, political, or social characteristics of real-world algorithmic deployments posed barriers for contestation, accountability, and abandonment? How can future calls for abandonment be bolstered and supported?
@@
In this paper, we take a first step toward answering these questions by exploring the concept of algorithm abandonment. We define algorithm abandonment as a decision made by actors with jurisdiction over the system, to discontinue the process of developing, deploying, or using the algorithm due to its (potential) harms. The decision to abandon an algorithm can be made at any point in the algorithm’s development lifecycle (Hub, 2022d; Suresh and Guttag, 2021), and includes both decisions not to build a proposed technology, and decisions to stop using tools that are already built. The algorithmic systems that we explore span diverse ways of coming into being, including hand-coded rules and machine learning models. Our study of algorithm abandonment draws from literature from Science and Technology Studies which examines algorithms not just as technical artifacts mapping inputs to outputs, but rather as socio-technical “assemblages” of social relationships that enable and sustain technologies (Schwennesen, 2019). Thus, understanding abandonment requires examining the social, political, and economic systems that algorithms are embedded within (Seaver, 2019).
@@
In our thematic analysis of 40 different campaigns to abandon harmful algorithms in recent years, we observe that the fall of an algorithm (or “dynamics” leading to abandonment) often follows a consistent pattern of six iterative, non-linear phases which we term the 6 “D”s of abandonment: (1) discovery, (2) diagnosis, (3) dissemination, (4) dialogue, (5) decision, and (6) death, which we visualize for three example cases in Figure 1. We review each of these phases and highlight both tactics that algorithm owners used to avoid scrutiny (e.g., discrediting or ignoring criticism), and tactics that other actors such as impacted individuals, advocacy organizations, journalists, or regulators used to shift power to advocates of abandonment. We additionally identify key socio-technical factors relevant to the outcome of whether an algorithm was ultimately abandoned (and how long it took to reach this stage), such as public awareness of and access to the algorithm, characteristics of the algorithm’s implementation, visibility of criticism, the regulatory environment, and for whom (and how) the algorithm is marketed. We conclude by discussing implications for technologists designing algorithmic systems, regulators and oversight entities, and the FAccT community in identifying barriers to abandonment and interventions to support future calls to abandon harmful algorithms. We release our annotated case set online111https://njohnson99.github.io/fall-of-algorithm-database/ as a resource to aid future research and advocacy efforts.
@@ 2. Background & Related Work
We contextualize our study of abandonment using scholarship that studies how different stakeholders respond to algorithmic harm. “Algorithmic harm,” as described by Shelby et al. (2023), refers to “the adverse lived experiences resulting from a system’s deployment and operation in the world” which occurs through the interplay of technical components and societal power dynamics (Noble, 2018; Benjamin, 2019; McGregor, 2020). We group approaches taken to address algorithmic harm into two broad categories: (1) actions taken “from above” by organizations in positions of power (e.g., those that develop, oversee, or govern the algorithm’s use), and (2) actions taken “from below” by other stakeholders (e.g., those impacted by algorithmic harm). We adopt the language of “above” and “below” from Whitney et al. (2021), who uses these terms to articulate the “tensions between the wants, needs, and knowledge of unequally empowered stakeholders”.
@@
We situate our study of abandonment as an organizational decision made “from above” by algorithm owners or government oversight entities. We use the term “algorithm owner” to broadly refer to the organizations and people within them that have the power to influence an algorithm’s (dis)continued use on the ground. Our analysis of the events, or “dynamics”, that precede abandonment in practice draws from the growing body of scholarship that examines how actors “from below” organize to contest harmful algorithmic systems (Zong and Matias, 2023; DeVrio et al., 2023). We use the term “impacted communities” to describe the people who are affected by the use of an algorithm. Those impacted may include the direct consumers of a technology or any other people who are subject to an algorithm’s decisions (Friedman et al., 2006; Moss et al., 2021).
@@
Repair from above. Much work concerning algorithmic harm explores the actions that algorithm owners might take to repair flawed technologies. The AI/ML and FAccT community have proposed numerous technical methods to promote values of “fairness” and “explainability” (Doshi-Velez and Kim, 2017; Selbst and Barocas, 2018; Liao and Vaughan, 2023). This work has led to the development of metrics (Speicher et al., 2018; Pessach and Shmueli, 2022; Barocas et al., 2023; Subhash et al., 2022), methods (Dwork et al., 2012; Wang and Rudin, 2015; Hardt et al., 2016; Ribeiro et al., 2016), and toolkits (Bellamy et al., 2018; Arya et al., 2019; Deng et al., 2022; Weerts et al., 2023) that practitioners can use to measure and fix specific harmful behaviors. Beyond technical repairs, human-computer interaction researchers have documented how organizational users and decision-makers engage in their own ad-hoc or improvisational repair activities when they encounter flawed technologies (Houston et al., 2016; Eslami et al., 2017; Schwennesen, 2019). For example, social workers using AI tools to screen calls can override them to mitigate disparities in final outcomes (De-Arteaga et al., 2020; Cheng and Chouldechova, 2022; Cheng et al., 2022). Such methods can be used to correct the algorithm’s flaws, often defined in terms of narrowly scoped measures (Jacobs and Wallach, 2021), while maintaining the algorithm’s operation.
@@
Refusal from below. The concept of refusal, defined simply as the act of saying “no” (Ahmed, 2017), has been studied in feminist and de-colonial discourses (Simpson, 2007; TallBear, 2013; McGranahan, 2016; Honig, 2021). Today, scholars and activists continue to draw from these traditions in studying refusal of emerging technologies (Cifor et al., 2019; DeVrio et al., 2023; Zong and Matias, 2023). Among theories of anti-techno-solutionism, Benjamin (2016) refers to indigenous refusal to analyze vignettes where research subjects refuse to engage with new biotechnologies. A group of scholars drafted the Feminist Data “Manifest-No”, a declaration of refusal to harmful data regimes and exploitative research practices (Cifor et al., 2019). There is a rich history of grassroots social movements formed to refuse technologies, tracing back hundreds of years to the British Luddites during the Industrial Revolution (Conniff, 2011), sparking modern adaptations of “Luddism” by workers organizing to oppose harmful algorithmic management practices (Chayka, 2023; Merchant, 2023).
Refusal of harmful technologies may take the form of individual users choosing to no longer interact with harmful algorithms (DeVrio et al., 2023; Zong and Matias, 2023). For example, in response to Amazon’s introduction of AI-powered cameras for worker surveillance, some drivers engaged in an act of refusal by obscuring the cameras with stickers, rendering them unable to record (Gurley, 2021). But refusal also extends beyond individual acts into collective calls for a system to be abandoned. Understanding calls to abandon an algorithm through the lens of refusal sheds light on the possibilities that abandonment entails to “open up possibilities for power shifts” (Casemajor et al., 2015).
@@ 3. Methodology
Case collection. To develop our dataset, we conducted an extensive search of various sources for instances where a harmful algorithm was either (a) already abandoned, or (b) critics called for the algorithm to be abandoned, regardless of whether it was abandoned or not at the time of writing. Our search methodology was a dynamic, iterative process that unfolded concurrently with our analysis. This approach involved consulting a diverse group of resources, detailed further below:
@@
AI incident databases: We reviewed all of the reported incidents to date222We conducted our review from July to December 2023, and reviewed all incidents that had been posted in each database before January 1, 2024. in three incident databases that compiled and documented real-world harms caused by algorithmic systems: (a) the AI Incident Database (AIID) (McGregor, 2020), (b) the AI, Algorithmic, and Automation Incidents and Controversies (AIAAIC) repository (AIAAIC, 2023), and (c) Not My AI (Varon and Peña, 2022) which focuses specifically on mapping cases of algorithmic harm in Latin America.
@@
International cases: Given that U.S. and European-based incidents were over-represented in our initial case set, we extended our search to be intentionally inclusive of other regions globally. We conducted additional keyword searches on Google to identify relevant cases that occurred in countries outside of Western contexts, such as Korea (Kwon and Yun, 2021), Argentina (Aplicada, 2018), Jordan (Ryan-Mosley, 2023a), India (amn, 2022), among others.
@@
| U.S. National Eating Disorder Assocation (NEDA)’s Tessa Wellness Bot (Wells, 2023) (chatbot) | Shortly after NEDA fired their hotline staff, activists posted on social media about how Tessa (NEDA’s chatbot) gave harmful and triggering advice. |
@@
From the cases that met the abandonment criteria, we selected a subset of 40 as our case studies for our thematic analysis. We chose this subset based on several criteria, including (a) coverage of different domains, (b) access to enough sources to gain more information about a case to do a thematic analysis, and (c) reaching saturation in finding new themes in the analysis process. The selected cases included algorithms used in social media, chatbots, hiring, family and social services, benefits allocation, healthcare, education, property tech, and law enforcement. We chose among cases within the same domain based on the amount of public information and documentation that was available for us to construct accurate event timelines. We performed our thematic analysis in parallel to identifying relevant cases, and stopped adding new cases to the subset once we had reached saturation in finding new themes (i.e., the types of events and factors relevant to abandonment). Our final case set contains 40 different algorithmic systems, 24 of which had been abandoned at the time of writing333Some algorithms that we study were deployed in multiple different contexts (e.g., PredPol was used by dozens of U.S. police departments (Haskins, 2019)), and were abandoned in some, but not others. We count products like PredPol as a single “algorithmic system”, and count them as being “abandoned” if they were abandoned in at least one such context. , while the remainder were still in use. 16 out of the 40 total cases involved campaigns that took place in countries outside of the U.S. To provide context, we describe three example cases in Table 1 and a complete list in Appendix C. These 40 cases are not intended to be a comprehensive list of all algorithms that were called to be abandoned, but rather serve as a starting point to conceptualize algorithm abandonment, illustrate its dynamics, and map out its implications.
@@
Thematic analysis. We gathered and reviewed public resources about each case, including news articles, research publications, social media posts, and any other relevant primary sources (e.g., algorithm source code or court documents). For each case, we constructed a timeline by noting the month and year of important events. Our goal in noting important events was not to be comprehensive, but rather to provide “good enough” coverage of the types of events leading to abandonment, a standard used in past qualitative research (Yildirim et al., 2023). We reviewed sources to code metadata about the algorithm, such as its application domain, the type(s) of algorithmic harm (Shelby et al., 2023), and other information relevant to whether it was ultimately abandoned (more details in Appendix A). We walk through the set of sources and resulting codes and event timeline for an example case in Appendix A.1.
@@
We conducted a bottom-up thematic analysis (Braun and Clarke, 2006) of case event timelines and metadata in a series of meetings among research team members. In our thematic analysis of event timelines, we paid particular attention to the sequencing of key events, the time lapsed between them, and how each event contributed to the final decision of whether the system was abandoned. Our analysis yielded six high-level types of events (or “phases”) as the dynamics of the algorithm abandonment that we describe in Section 4. To identify factors relevant to abandonment presented in Section 5, we noted differences among algorithms that were abandoned almost immediately (e.g., within days of their release), versus algorithms that took months or even years to be abandoned. We organized these differences into higher-level categories that resulted in the final seven factors. We also noted different barriers that advocates of abandonment faced when interacting with algorithm owners (i.e., asserting their role in this accountability relationship (Metcalf et al., 2023)).